# Exploring Neural Granger Causality with xLSTMs: Unveiling Temporal Dependencies in Complex Data

### v4

@@ 3.3 Theoretical Analysis
Ultimately, we want to ensure that, provided real-world data, we can find hyperparameters such that Algorithm 1 discovers all and only those edges of the unique underlying GC graph $(\mathcal{V},\mathcal{E})$ as per Definition 1. Providing convergence guarantees in full generality is notoriously hard for such practical architectures and optimization schemes, and thus rarely attempted. However, we can at least investigate whether the chosen model class containing $\mathcal{M}_{\bm{\theta},v}$ can approximate $g_{v}$ to arbitrary precision. If that is the case, we can be reasonably certain that gradient-based optimization schemes will yield satisfactory approximations, even in the absence of formal guarantees.
The forecasting component of GC-xLSTM consists of two main steps, namely the sparse initial projection $\bm{W}_{v}$ and the subsequent xLSTM blocks. One might think that the sparsity of $\bm{W}_{v}$ hinders learning the correct $g_{v}$. However, the true underlying $g_{v}$ is independent of all variates $w$ without ingoing edges into variate $v$. Thus, depending on an appropriate choice of the sparsity hyperparameter $\lambda$, the projection $\bm{W}_{v}$ can encode exactly those as zero entries. It remains to investigate the approximation capabilities of the sLSTM blocks, which we present in Appendix B in more detail. In summary, we can assume sLSTM blocks to be at least as powerful as RNNs, which are, themselves, universal function approximators. Thus, the overall GC-xLSTM architecture is sufficiently rich to model $g_{v}$ to adequate precision. We continue by confirming this empirically in the next section.
@@ 4 Experimental Evaluation
Table 1: GC-xLSTM is highly accurate at discovering GC relations in the chaotic and non-linear Lorentz-96 system. For each setting and baseline, we provide the accuracy, balanced accuracy (BA), and AUROC. The best models are highlighted bold.
@@
Model Accuracy (↑) BA (↑) AUROC (↑) Accuracy (↑) BA (↑) AUROC (↑)
VAR 91.8
@@
We conduct extensive experiments on six datasets to assess the practical effectiveness of GC-xLSTM. We will now explain the chosen parameters used to train the GC-xLSTM architecture and then discuss the used datasets before presenting our obtained results in detail.
@@
Architecture Details.   For our component-wise networks, we use a single xLSTM block comprised of one sLSTM layer, followed by a linear layer to predict the next $H=10$ from the preceding $C=10$ time steps (for the VAR dataset, we set both to $5$). The hidden dimension of the sLSTM block is set to 32 for all datasets. We find that the presence of a gated MLP to up- and down-project the hidden states of the sLSTM block does not significantly improve performance, so we omit it in all our experiments for simplicity. We deliberately do not use any mLSTM blocks, as we find that the sLSTM blocks are superior at capturing long-range dependencies in the data (cf. Kraus et al. [2024]).
@@
Evaluation Details.   We use the Adam optimizer [Kingma and Ba, 2017] for full gradient descent training with a weight decay of 0.1. We schedule the learning rate to follow a linear warmup of 2,000 iterations to $\eta=10^{-4}$, followed by cosine annealing until the end of training for a total of 13,000 steps. We start learning the reduction coefficients after a warmup period of $K=1,500$ iterations, during which the uniform compression across all columns combined with the prediction loss gives reasonable priors for the gradient directions of the reduction coefficients. Due to the moderate size of the datasets, we performed full-batch gradient descent. Only the sparsity hyperparameter $\lambda$ was tuned specifically for each of the settings. This allows obtaining degrees of sparsity specifically tailored to the characteristics and requirements of each dataset and task. Note that, except for the customization of $\lambda$, we use essentially the same hyperparameter configuration for a very wide set of datasets, underpinning the robustness of GC-xLSTM. Following [Tank et al., 2022, Sec. 6.1], we compute AUROC scores by sweeping over $\lambda\in\{5,\dots,15\}$. All training runs were carried out on a single NVIDIA RTX A6000 GPU and concluded in at most 1.5 hours.
@@
Datasets.   We evaluate GC detection with GC-xLSTM on six diverse datasets. Obtaining objective truth about the underlying graph for real-world scenarios is a constant challenge in Granger causality research. We thus employ the Lorenz-96 system of differential equations [Karimi and Paul, 2010], realistic fMRI brain activity simulations [Smith et al., 2011], and simulated linear VAR data following Tank et al. [2022]. For further qualitative insights on real-world data, we additionally analyze the Moléne weather dataset [Girault, 2015], human motion capture recordings [CMU, 2009], and company fundamentals [Divo et al., 2024]. More details on the data sets are provided in Appendix C.
@@
Figure 3: Granger causal relations on the Moléne Dataset. GC-xLSTM uncovers several interesting Granger causal relations giving us an insight into the dynamic weather patterns. We can observe that the sparsity of the learned Granger causal relations increases with the value of $\lambda$.
@@ 4.1 Main Results
Lorenz-96.   Table 1 evaluates GC detection with GC-xLSTM on three metrics. As a qualitative comparison, we provide results for seven well-known baselines: VAR as classic $F$-tests for Granger causality taken from Marcinkevičs and Vogt [2021], cMLP and cLSTM [Tank et al., 2022], GC-KAN [Lin et al., 2024], TCDF [Nauta et al., 2019], eSRU [Khanna and Tan, 2020], and GVAR [Marcinkevičs and Vogt, 2021]. GC-xLSTM outperforms all baseline methods for both $F=10$ and $F=40$ in two out of three metrics. This shows that GC-xLSTM captures the underlying Granger causal relationships in the presence of limited and noisy data. Furthermore, the model displays strong prediction accuracies as Figure 5 confirms visually.
@@
fMRI.   Next, we evaluated the efficacy of GC-xLSTM in noisy settings by considering the rich and realistic simulations of BOLD deconvolved time series provided by the fMRI data. As the balanced accuracies (BA) in Table 2 show, GC-xLSTM significantly outperforms the baseline models.
Moléne.   Unlike Zoroddu et al. [2024], who incorporate graph prior knowledge based on sensor locations, our approach learns the GC structure solely from the temperature observations. This ensures that GC-xLSTM does not inherently favor regional connections over long-range dependencies, allowing it to discover dominant weather patterns operating both locally and across broader spatial scales. Adjusting $\lambda$ allows balancing granularity and interpretability for insights into both local and regional dependencies. The dense structure of the resulting Figure 3(a) exhibits a richer set of GC interactions, while the more sparse Figure 3(b) highlights only the most pronounced edges.
Human Motion Capture.   To analyze GC relations between body joints, we focus on two specific activities: Salsa dancing and running, which provide interpretable motion patterns. Figure 4 shows the results of the learned graphs for those activities. A closer look offers an intuitive understanding of the learned interactions. For example, in Salsa dance, we observe edges from the feet to the knees and the arms, supporting the characteristic movements of the lower driving the upper body. We can also see the cross-limb correlation, with movements initiating on one side of the body and being propagated to the other. Similarly, the results for running strongly establish the lower limbs as primary motion drivers, with edges from the feet and knees to the arms. We also see cyclic dependencies between the knees, ankles, and feet, capturing the repetitive, alternating nature of the gait.
Figure 6: GC-xLSTM extracs relations between company fundamentals.
Company Fundamentals.   Lastly, we evaluated how well GC-xLSTM can uncover relationships between financial indicators of large companies. As the excerpt in Figure 6 shows and as verified by a financial expert, the extracted edges are mostly economically sensible. Full results are provided in Appendix D.
@@ 4.2 Model Analysis
Inspecting Training Dynamics.   Here, we elaborate on the utility of the logarithm in the reduction loss $\mathcal{L}_{\text{red}}$ of Eq. 3. The logarithm term incentivizes the model to explore sparser solutions to GC discovery by allowing the training to move forward over any local minima that make use of the full set of input variates. As the projection matrix becomes sparser and the input variates vanish from consideration by the model, the prediction loss $\mathcal{L}_{\text{pred}}$ will slightly increase. As seen in Figure 8 in Appendix E, an increase in the sparsity of the feature selectors $\bm{W}$ drives down the loss and enables the model to learn more meaningful GC relations. It also shows how the variable usage quickly stabilizes.
@@
Variable Lags.   GC-xLSTM can naturally be extended to learn separate projections $\bm{W}^{(\ell)}$ per time lag $\ell$, effectively inflating $\bm{W}$ it to rank three. We evaluate it on the simulated VAR dataset across different time series lengths $T$ and number of variates $V$ in Table 3, proving its ability to learn multiple lag-specific relations without training additional xLSTM models.
Complexity and Scaling Behaviour.   To discover the entire GC graph, we need to fit $V$ models $\mathcal{M}_{\bm{\theta}_{v},v}$ to obtain the respective sparse projection matrices $\bm{W}_{v}$ containing all edges arriving at each $v$. Assuming for brevity that the latent dimension $D\propto V$, i.e., is a fixed multiple of the number of variates $V$, each fitting runs in $O(TV^{2}/N_{h})$ time and requires $O(TV^{2}/N_{h})$ space, with $T$ being the time steps and $N_{h}$ the number of sLSTM heads. Their block-wise structure permits that, despite the squared effort, forward and backward passes are efficiently computable even for thousands of dimensions. Depending on whether all $V$ models are estimated in sequence or parallel, either the time or memory complexity multiplies by $V$ to arrive at the total cost. In practice, however, GC-xLSTM is extremely efficient on contemporary computing platforms due to the availability of highly optimized implementations for xLSTM layers. Figure 9 in Appendix F shows that it effectively scales linearly in the number of variates for the ranges relevant to common GC detection settings.
@@ 5 Conclusion
Limitations.   While Section 3.3 provides a theoretical analysis of GC-xLSTM, it does not provide guarantees in the form of mathematical proofs. The rigor is limited by a lack of formal analysis of xLSTM blocks (see also Appendix B), which are yet to be formally analyzed to the same degree as more established architectures like LSTMs. While we discuss and measure the scaling behaviour of GC-xLSTM in Section 4.2, we only focused on dozens of variates as common in the literature.
Future work.   This includes using more sophisticated architectures such as xLSTM-Mixer [Kraus et al., 2024] or TimeMixer [Wang et al., 2024]. Furthermore, the discovery of causal links specific to certain lags could be refined, where per-lag projections are learned for the near past and a remainder projection for the more distant lags. Finally, extending our evaluations to more real-world datasets encompassing domains such as climate change or ecology is an important next step.

### v5

@@ 3.3 Theoretical Analysis
Ultimately, we want to ensure that, provided real-world data, we can find hyperparameters such that Algorithm˜1 discovers all and only those edges of the unique underlying GC graph $(\mathcal{V},\mathcal{E})$ as per Definition˜1. Providing convergence guarantees in full generality is notoriously hard for such practical architectures and optimization schemes, and thus rarely attempted. However, we can at least investigate whether the chosen model class containing $\mathcal{M}_{\bm{\theta},v}$ can approximate $g_{v}$ to arbitrary precision. If that is the case, we can be reasonably sure that gradient-based optimization schemes will yield satisfactory approximations, even without formal guarantees.
The forecasting component of GC-xLSTM consists of two main steps: the sparse initial projection $\bm{W}_{v}$ and the subsequent xLSTM blocks. One might think that the sparsity of $\bm{W}_{v}$ hinders learning the correct $g_{v}$. However, the true underlying $g_{v}$ is independent of all variates $w$ without ingoing edges into variate $v$. Thus, depending on an appropriate choice of the sparsity hyperparameter $\lambda$, the projection $\bm{W}_{v}$ can encode exactly those as zero entries. It remains to investigate the approximation capabilities of the sLSTM blocks, which we present in Appendix˜B in more detail. In summary, we can assume that sLSTM blocks are at least as powerful as RNNs, which are, in turn, universal function approximators. Thus, the overall GC-xLSTM architecture is sufficiently rich to model $g_{v}$ to adequate precision. We continue by confirming this empirically in the next section.
@@ 4 Experimental Evaluation
Table 1: GC-xLSTM is highly accurate at discovering GC relations in the chaotic and non-linear Lorentz-96 system. For each setting and baseline, we provide the accuracy (Acc.), balanced accuracy (BA), and AUROC. The best models are highlighted as bold.
@@
Model Year Acc. (↑) BA (↑) AUROC (↑) Acc. (↑) BA (↑) AUROC (↑)
VAR 2021 91.8
@@
We conduct extensive experiments on six datasets to assess the practical effectiveness of GC-xLSTM. We will now explain the chosen architecture and parameters used to train GC-xLSTM and then discuss the datasets used before presenting our obtained results in detail. The results are split into investigating the general GC modeling capabilities on a diverse range of applications (Section˜4.1) and a subsequent model analysis, including an ablation study and using different numbers of variates (Section˜4.2).
Architecture Details.
@@
For our component-wise networks, we use a single xLSTM block comprised of one sLSTM layer, followed by a linear layer to predict the next time step from the preceding $C=10$ time steps (see Table˜5 in Appendix˜C for deviations). We then directly optimize the architecture on each time step. The hidden dimension of the sLSTM block is set to 32 for all datasets. We find that the presence of a gated MLP to up- and down-project the hidden states of the sLSTM block does not significantly improve performance, so we omit it in all our experiments for simplicity. We deliberately do not use any mLSTM blocks, as we also find that the sLSTM blocks are superior at capturing long-range dependencies in the data [Kraus et al., 2025].
Training and Evaluation Details.
@@
We use the Adam optimizer [Kingma and Ba, 2017] for full gradient descent training with a weight decay of 0.1. We schedule the learning rate to follow a linear warmup of 2,000 iterations to $\eta=10^{-4}$, followed by cosine annealing until the end of training for a total of 13,000 steps. We start learning the reduction coefficients after a warmup period of $K=1,500$ iterations, during which the uniform compression across all columns combined with the prediction loss gives reasonable priors for the gradient directions of the reduction coefficients. Due to the moderate size of the datasets, we performed full-batch gradient descent. Only the sparsity hyperparameter $\lambda$ was tuned specifically for each setting. This allows obtaining degrees of sparsity specifically tailored to the characteristics and requirements of each dataset and task. Note that, except for the customization of $\lambda$, we use essentially the same hyperparameter configuration for an extensive set of datasets, underpinning the robustness of GC-xLSTM. Following [Tank et al., 2022, Sec. 6.1], we compute all AUROC scores by sweeping over $\lambda\in\{5,\dots,15\}$ in steps of one. This effectively integrates this hyperparameter as it covers the entire empirically viable range. All training runs were carried out on a single NVIDIA RTX A6000 GPU and concluded in at most 1.5 hours. Mini-batching could further decrease this modest training time.
Datasets.
@@
We evaluate GC detection with GC-xLSTM on six diverse datasets. Obtaining objective truth about the underlying graph for real-world scenarios is a constant challenge in Granger causality research. We thus employ the Lorenz-96 system of differential equations [Karimi and Paul, 2010], realistic fMRI brain activity simulations [Smith et al., 2011], and simulated linear VAR data following Tank et al. [2022]. For further qualitative insights on real-world data, we additionally analyze the Moléne weather dataset [Girault, 2015], human motion capture recordings [CMU, 2009], and company fundamentals [Divo et al., 2025]. An overview and more details are provided in Appendix˜C.
@@
Figure 3: GC-xLSTM uncovers dynamic GC weather patterns in the Moléne dataset. We observe that the sparsity of the learned Granger causal relations increases with higher $\lambda$.
@@ 4.1 Main Results
Lorenz-96.
Table˜1 evaluates GC detection with GC-xLSTM on three metrics. As a qualitative comparison, we provide results for seven well-known baselines: VAR as classic $F$-tests for Granger causality taken from Marcinkevičs and Vogt [2021], cMLP and cLSTM [Tank et al., 2022], GC-KAN [Lin et al., 2024], TCDF [Nauta et al., 2019], eSRU [Khanna and Tan, 2020], and GVAR [Marcinkevičs and Vogt, 2021]. Their scores are taken as reported in the original papers. First, GC-xLSTM outperforms all baseline methods for both $F=10$ and $F=40$ in both accuracy and balanced accuracy. This shows that GC-xLSTM reliably captures the underlying Granger causal relationships in the presence of limited and noisy data. Second, it demonstrates that the sparsity hyperparameter $\lambda$ is rather well-behaved, as the AUROC score resulting from a sweep over a range of $\lambda$ values still provides very competitive results. Third, as a qualitative validation, Figure˜5 visually confirms the strong prediction accuracies of GC-xLSTM.
Moléne.
Unlike Zoroddu et al. [2024], who incorporate graph prior knowledge based on sensor locations, our approach learns the GC structure solely from the temperature observations. This ensures that GC-xLSTM does not inherently favor regional connections over long-range dependencies, allowing it to discover dominant weather patterns operating both locally and across broader spatial scales. Adjusting $\lambda$ allows balancing granularity and interpretability for insights into both local and regional dependencies. The dense structure of the resulting Figure˜3(a) exhibits a richer set of GC interactions, while the more sparse Figure˜3(b) highlights only the most pronounced edges.
@@
fMRI.
Next, we evaluated the efficacy of GC-xLSTM in noisy settings by considering the rich and realistic BOLD deconvolved time series provided by the fMRI data. As the balanced accuracies (BA) in Table˜2 show, GC-xLSTM significantly outperforms the baseline models.
Human Motion Capture.
To analyze GC relations between body joints, we focus on two specific activities: Salsa dancing and running, which provide interpretable motion patterns. Figure˜4 shows the results of the learned graphs for those activities. A closer look offers an intuitive understanding of the learned interactions. For example, in the Salsa dance, we observe edges from the feet to the knees and the arms, supporting the characteristic movements of the lower driving the upper body. We can also see the cross-limb correlation, with movements initiating on one side of the body and propagating across. Similarly, the results for running strongly establish the lower limbs as primary motion drivers, with edges from the feet and knees to the arms. The cyclic dependencies between the knees, ankles, and feet capture the repetitive, alternating nature of the gait.
Figure 6: GC-xLSTM extracts relations between company fundamentals.
Company Fundamentals.
Lastly, we evaluated how well GC-xLSTM can uncover relationships between financial indicators of large companies. The dataset consists of short time series of quarterly performance indicators of 2527 large publicly traded companies. As the excerpt in Figure˜6 shows and as verified by a financial expert, those extracted edges are mostly economically sensible. Results on the full set of 19 features are provided in Appendix˜D.
@@ 4.2 Model Analysis
Ablation Study.
GC-xLSTM comprises two key innovations: Employing the xLSTM architecture (cf. Section˜3.1) and a novel joint optimization strategy (cf. Section˜3.2). To disentangle the empirical impact of each, we performed an ablation study. Table˜3 to the right shows balanced accuracy results on the Lorentz ($F=40$) and fMRI datasets.
+0.5 Ablation Forecaster Optimization Lorentz fMRI GC-xLSTM xLSTM Joint 96.6 ±0.3 73.3 ±3.0 (I) xLSTM Joint 93.0 ±0.3 62.8 ±2.0 (II) xLSTM Group Lasso 73.0 ±4.6 65.4 ±2.0
Table 3: Both the xLSTM forecaster and the novel joint optimization of GC-xLSTM drive its improved performance.
First, to assess the architectural contribution, we replaced the xLSTM block with a standard LSTM. The resulting substantial performance drop (GC-xLSTM → (I)) validates the importance of the xLSTM’s advanced modeling capacity. Second, to evaluate the optimization strategy, we substituted it with standard Group Lasso [Simon and Tibshirani, 2012]. This also led to a marked decline in performance (GC-xLSTM → (II)), confirming its ability to enforce strict input sparsity.
@@
Variable Number of Lags.
GC-xLSTM can naturally be extended to learn separate projections $\bm{W}^{(\ell)}$ per time lag $\ell$, effectively inflating $\bm{W}$ to rank three. We evaluate it on the simulated VAR dataset across different time series lengths $T$ and number of variates $V$ in Table˜4, proving its ability to learn multiple lag-specific relations without training additional xLSTM models.
Complexity and Scaling Behaviour.   To discover the entire GC graph, we need to fit $V$ models $\mathcal{M}_{\bm{\theta}_{v},v}$ to obtain the respective sparse projection matrices $\bm{W}_{v}$ containing all edges arriving at each $v$. Assuming for brevity that the latent dimension $D\propto V$, i.e., is a fixed multiple of the number of variates $V$, each fitting runs in $O(TV^{2}/N_{h})$ time and requires $O(TV^{2}/N_{h})$ space, with $T$ being the time steps and $N_{h}$ the number of sLSTM heads. Their block-wise structure permits that, despite the squared effort, forward and backward passes are efficiently computable even for thousands of dimensions. Depending on whether all $V$ models are estimated in sequence or parallel, either the time or memory complexity multiplies by $V$ to arrive at the total cost. In practice, however, GC-xLSTM is extremely efficient on contemporary computing platforms due to the availability of highly optimized implementations for xLSTM layers. Figure˜8 in Appendix˜E shows that it effectively scales linearly in the number of variates for the ranges relevant to standard GC detection settings.
Inspecting Training Dynamics.   Here, we elaborate on the utility of the logarithm in the reduction loss $\mathcal{L}_{\text{red}}$ of Eq.˜3. The logarithm term incentivizes the model to explore sparser solutions to GC discovery by allowing the training to move forward over any local minima that use the complete set of input variates. As the projection matrix becomes sparser and the input variates vanish from consideration by the model, the prediction loss $\mathcal{L}_{\text{pred}}$ will slightly increase. As seen in Figure˜9 in Appendix˜F, an increase in the sparsity of the feature selectors $\bm{W}$ drives down the loss and enables learning more meaningful GC relations. It also shows how the variable usage quickly stabilizes.
@@ 5 Conclusion
Limitations.   While Section˜3.3 provides a theoretical analysis of GC-xLSTM, it does not give guarantees in the form of mathematical proofs. The rigor is limited by a lack of formal analysis of xLSTM blocks (see also Appendix˜B), which are yet to be formally analyzed to the same degree as more established architectures like LSTMs. While we discuss and measure the scaling behaviour of GC-xLSTM in Section˜4.2, we only focused on dozens of variates as common in the literature.
Future work.   This includes using more sophisticated architectures such as TimeMixer [Wang et al., 2024] or xLSTM-Mixer [Kraus et al., 2025]. Furthermore, discovering causal links specific to certain lags could be refined, where per-lag projections are learned for the near past and a remainder projection for the more distant lags. Finally, extending our evaluations to more real-world datasets encompassing domains such as climate change or ecology is an essential next step.